Thermodynamics

Q2

using the following data, calculate the enthalpy of solution chloride:


Enthalpy of hydration of Li+ = -520 kJ mol-1

Enthalpy of hydration of CI- = -364 kJmol -1

Lattice enthalpy LiCI = -860 kJ mol-1


(2 marks)

please detailed and correct answers 🙏🏾🙏🏾


1
Expert's answer
2022-01-18T14:10:02-0500

Enthalpy is the sum of the systems internal energy and the product of its pressure and volume.

Enthalpy= (h2-h1)=cp( T2-T1 )